# Where Do People Tell Stories Online? Story Detection Across Online Communities

### v2

@@ A.4 In what conversational contexts do people tell stories?
| r/askscience |
6.35 |
0.56 |
0.09 |
| r/philosophy |
3.83 |
0.36 |
0.10 |
| r/legaladvice |
3.58 |
0.97 |
0.27 |
| r/NoStupidQuestions |
3.47 |
0.68 |
0.19 |
| r/summonerschool |
3.40 |
0.62 |
0.18 |
| r/LeagueofLegendsMeta |
3.38 |
0.41 |
0.12 |
| r/Bitcoin |
3.08 |
0.53 |
0.17 |
| r/applehelp |
3.07 |
0.77 |
0.25 |
| r/techsupport |
3.02 |
0.82 |
0.27 |
| r/poker |
2.88 |
0.77 |
0.27 |
@@
| r/nfl |
1.21 |
0.50 |
0.41 |
| r/LifeProTips |
1.21 |
0.62 |
0.52 |
| r/weddingplanning |
1.20 |
0.92 |
0.77 |
| r/TalesFromRetail |
1.19 |
0.97 |
0.82 |
| r/DoesAnybodyElse |
1.16 |
0.77 |
0.66 |
| r/travel |
1.16 |
0.76 |
0.66 |
| r/harrypotter |
1.16 |
0.81 |
0.70 |
| r/AskReddit |
1.10 |
0.88 |
0.80 |
| r/SquaredCircle |
1.10 |
0.70 |
0.64 |
| r/Random_Acts_Of_Amazon |
1.01 |
0.90 |
0.89 |
@@ A.6 Prior Definitions of Stories
world building or setting |
Piper et al. (2021a) |
@@
readability |
dos Santos et al. (2017) |
verb classes |
Eisenberg and Finlayson (2017) |
syntactic production rules |
@@
verb sequence perplexity |
Yao and Huang (2018) |
dependency tags |
Piper et al. (2021a) |
tense |
Piper et al. (2021a) |
mood |
Piper et al. (2021a) |
voice |
Piper et al. (2021a) |
amount of dialog |
Piper et al. (2021a) |
named entities |
Ceran et al. (2012) |
stative verbs |
Ceran et al. (2012) |
semantic triples |
Ceran et al. (2012) |
@@ A.8.1 Few-Shot Document Classification
We modify the below prompt based on the setting. For zero-shot, we exclude the "Examples" section. We test two few-shot settings: k=2 and k=4 (where k is the number of examples). In each case, we sample from the training dataset until we have a total of k examples, evenly split between positive (i.e. story) and negative (i.e. not story) instances from the training dataset. We then interleave the negative and positive samples in the "Examples:" section.
We also try excluding the Guidelines section. We find that including the Guidelines section improves performance in the zero-shot setting. In contrast, the best performing few-shot setting for each model used k=4 examples and excluded the Guidelines section.
@@
A story describes a sequence of events involving one or more people.
@@
Answer: <YES/NO BASED ON GOLD_CONSENSUS>
Task:
@@ A.8.2 Chain-of-Thought Document Classification
Using the definitions for ’people’ and ’events’ given in the guidelines above, answer the following questions
Question 1: Does the text contain ’people’? If so, list them.
Question 2: Does the text contain a sequence of causally connected ’events’. If so, list them.
Finally, based on the guidelines and your answers to Question 1 and Question 2, decide whether the text contains a story. Respond ’Yes’ or ’No’.
@@
Example 1:
’Yesterday, I went to the store to buy a jug of milk so that I could make pancakes. When I arrived, the cashier told me that the store lost power the previous night, so all the milk spoiled. So much for pancakes!’
Question 1 Answers: [’I’, ’cashier’].
Question 2 Answers: [’went’, ’arrived’, ’told’, ’lost’, ’spoiled’].
Story Decision: Yes
Example 2:
’Not a good FDIC tip. Coverage levels start at $250k and can be increased by certain multipliers–so that’s not really a concern unless you’ve got several million in the bank.’ ’
Question 1 Answers: [].
@@
Your turn. Text:
<TEXT>
@@ A.8.3 Few-Shot Story Boundary Detection
Formulating a story boundary detection task for GPT-4 is difficult, due the proclivity for responses to include unrequested mutations to the original source string, which can lead to token misalignment. We mitigate this issue through careful prompt engineering using the train split of the dataset, use of manually-constructed few-shot examples that demonstrate desired behavior on different formats observed in our training data, and setting the temperature to 0 in OpenAI chat completion requests. The results support the token-level story span detection results (see Table 4). We include the prompt below. Finally, we note that in consideration of the poor performance of this method relative to the Fine-tuned RoBERTa model, in general we do not recommend this approach for for token-level discourse boundary detection tasks.
@@
Example 3: <b>Title:</b> I [24F] am about to fail my Bio class… again. Help. (<br><br><b>Post:</b> Idk what to do AGGHHHH!!!!?$$@#. Help please. I started studying religiously two weeks ago, and I already covered 4 chapters, so maybe I’m on the right track? I talked to the professor, but he didn’t offer much help. One TA sent me a list of additional resources, but it’s not realistic for me to read all of those in time for the exam.
Example 3 Answer: <b>Title:</b> I [24F] am about to fail my Bio class… again. Help. (<br><br><b>Post:</b> Idk what to do AGGHHHH!!!!?$$@#. Help please. <<S>>I started studying religiously two weeks ago, and I already covered 4 chapters, so maybe I’m on the right track? I talked to the professor, but he didn’t offer much help. One TA sent me a list of additional resources, but it’s not realistic for me to read all of those in time for the exam.<<E>>
@@ A.9.2 Story Feature Comparison Across Datasets
In section 6.1, we conduct -tests comparing features between texts labeled as containing stories vs. not containing stories in the
@@
| Measure | Effect Size () | Direction |
-value |
Effect Size () | Direction |
-value |
@@
expert-annotated events |
1.899*** |
story |
n/a |
n/a |
n/a |
realis events |
1.429*** |
story |
1.687*** |
story | ||
past tense |
1.408*** |
story |
1.207*** |
story | ||
1st-person singular pronouns |
1.009*** |
story |
0.84*** |
story | ||
concreteness |
0.439*** |
story |
1.526*** |
story | ||
3rd-person singular pronouns |
0.397*** |
story |
1.281*** |
story | ||
entity mentions |
0.285** |
story |
0.006 |
1.085*** |
story | |
non-past tense |
0.947*** |
non-story |
0.639 |
is comment (vs. post) |
0.612*** |
non-story |
n/a |
n/a |
n/a |
2nd-person pronouns |
0.444*** |
non-story |
0.285* |
story |
0.017 |
sentence length |
0.259* |
non-story |
0.012 |
0.828*** |
non-story | |
| 1st-person plural pronouns |
0.106 |
0.154 |
| text length |
0.106 |
0.825*** |
non-story | |

### v3

@@ A.4 In what conversational contexts do people tell stories?
| r/askscience | 6.35 | 0.56 | 0.09 |
| r/philosophy | 3.83 | 0.36 | 0.10 |
| r/legaladvice | 3.58 | 0.97 | 0.27 |
| r/NoStupidQuestions | 3.47 | 0.68 | 0.19 |
| r/summonerschool | 3.40 | 0.62 | 0.18 |
| r/LeagueofLegendsMeta | 3.38 | 0.41 | 0.12 |
| r/Bitcoin | 3.08 | 0.53 | 0.17 |
| r/applehelp | 3.07 | 0.77 | 0.25 |
| r/techsupport | 3.02 | 0.82 | 0.27 |
| r/poker | 2.88 | 0.77 | 0.27 |
@@
| r/nfl | 1.21 | 0.50 | 0.41 |
| r/LifeProTips | 1.21 | 0.62 | 0.52 |
| r/weddingplanning | 1.20 | 0.92 | 0.77 |
| r/TalesFromRetail | 1.19 | 0.97 | 0.82 |
| r/DoesAnybodyElse | 1.16 | 0.77 | 0.66 |
| r/travel | 1.16 | 0.76 | 0.66 |
| r/harrypotter | 1.16 | 0.81 | 0.70 |
| r/AskReddit | 1.10 | 0.88 | 0.80 |
| r/SquaredCircle | 1.10 | 0.70 | 0.64 |
| r/Random_Acts_Of_Amazon | 1.01 | 0.90 | 0.89 |
@@ A.6 Prior Definitions of Stories
“A narrative is a discourse presenting a coherent sequence of events which are causally related and purposely related, concern specific characters and times, and overall displays a level of organization beyond the commonsense coherence of the events themselves, such as that provided by a climax or other plot structure.” – Eisenberg and Finlayson (2021)
@@
“Operationalizing Smith (2001)’s characteristics, our codebook had four inclusion criteria: the presence of a plot, characters, the author as a character, and a clear beginning, middle,and end. Additionally, we included three exclusion criteria. Posts were marked as non-narrative that were: purely informational, entirely providing resources, or entirely composed of a question posed to the subreddit community.” – Doyle et al. (2024)
“A STORY follows a plot-like structure (e.g. has an introduction, middle section or conclusion) or contains a sequence of events.” – Falk and Lapesa (2023)
@@
Roos and Reccius (2021) Castricato et al. (2021) |
| world building or setting | Piper et al. (2021a) |
@@
dos Santos et al. (2017)
@@
| readability | dos Santos et al. (2017) |
| verb classes | Eisenberg and Finlayson (2017) |
| syntactic production rules | Yao and Huang (2018) |
| verb sequence perplexity | Yao and Huang (2018) |
| dependency tags | Piper et al. (2021a) |
| tense | Piper et al. (2021a) |
| mood | Piper et al. (2021a) |
| voice | Piper et al. (2021a) |
| amount of dialog | Piper et al. (2021a) |
| named entities | Ceran et al. (2012) |
| stative verbs | Ceran et al. (2012) |
| semantic triples | Ceran et al. (2012) |
@@ A.8.1 Few-Shot Document Classification
We modify the below prompt based on the setting. For zero-shot, we exclude the “Examples” section. We test two few-shot settings: and (where is the number of examples). In each case, we sample from the training dataset until we have a total of examples, evenly split between positive (i.e. story) and negative (i.e. not story) instances from the training dataset. Because we use k-fold cross validation, we sample examples from the fold-specific training dataset. We then interleave the negative and positive samples in the “Examples” section.
We also try excluding the Guidelines section. We find that including the Guidelines section improves performance in the zero-shot setting. In contrast, the best performing few-shot setting for each model used examples and excluded the Guidelines section.
@@
Use the guidelines below to support your decisions, but ultimately, follow your best judgment as there are many edge cases.A story describes a sequence of events involving one or more people.
@@
Answer: <YES/NO>Task:
@@ A.8.2 Chain-of-Thought Document Classification
Using the definitions for ‘people’ and ‘events’ given in the guidelines above, answer the following questions
Question 1: Does the text contain ‘people’? If so, list them.
Question 2: Does the text contain a sequence of causally connected ‘events’. If so, list them.
Finally, based on the guidelines and your answers to Question 1 and Question 2, decide whether the text contains a story. Respond ‘Yes’ or ‘No’.
@@
Example 1: ‘Yesterday, I went to the store to buy a jug of milk so that I could make pancakes. When I arrived, the cashier told me that the store lost power the previous night, so all the milk spoiled. So much for pancakes!’
Question 1 Answers: [‘I’, ‘cashier’].
Question 2 Answers: [‘went’, ‘arrived’, ‘told’, ‘lost’, ‘spoiled’].
Story Decision: YesExample 2:
‘Not a good FDIC tip. Coverage levels start at $250k and can be increased by certain multipliers--so that’s not really a concern unless you’ve got several million in the bank.’Question 1 Answers: [].
@@
Your turn.
Text: <TEXT>
@@ A.8.3 Few-Shot Story Boundary Detection
Formulating a story boundary detection task for GPT-4 is difficult, due the proclivity for responses to include unrequested mutations to the original source string, which can lead to token misalignment. We mitigate this issue through careful prompt engineering, including manually-constructed few-shot examples that demonstrate desired behavior on different formats observed in our data, and setting the temperature to 0 in OpenAI chat completion requests. The results support the token-level story span detection results (see Table 4). We include the prompt below. Finally, we note that in consideration of the poor performance of this method relative to the Fine-tuned RoBERTa model, in general we do not recommend this approach for for token-level discourse boundary detection tasks.
@@
Example 3: <b>Title:</b> I [24F] am about to fail my Bio class... again. Help. (<br><br><b>Post:</b> Idk what to do AGGHHHH!!!!?$$@#. Help please. I started studying religiously two weeks ago, and I already covered 4 chapters, so maybe I’m on the right track? I talked to the professor, but he didn’t offer much help. One TA sent me a list of additional resources, but it’s not realistic for me to read all of those in time for the exam.
Example 3 Answer: <b>Title:</b> I [24F] am about to fail my Bio class... again. Help. (<br><br><b>Post:</b> Idk what to do AGGHHHH!!!!?$$@#. Help please. <<S>>I started studying religiously two weeks ago, and I already covered 4 chapters, so maybe I’m on the right track? I talked to the professor, but he didn’t offer much help. One TA sent me a list of additional resources, but it’s not realistic for me to read all of those in time for the exam.<<E>>
@@ A.9.2 Story Feature Comparison Across Datasets
In §6.1, we conduct -tests comparing features between texts labeled as containing stories vs. not containing stories in the
@@
| Measure | Effect Size () | Direction | -value | Effect Size () | Direction | -value |
@@
| expert-annotated events | 1.899*** | story | n/a | n/a | n/a | |
| realis events | 1.429*** | story | 1.687*** | story | ||
| past tense | 1.408*** | story | 1.207*** | story | ||
| 1st-person singular pronouns | 1.009*** | story | 0.84*** | story | ||
| concreteness | 0.439*** | story | 1.526*** | story | ||
| 3rd-person singular pronouns | 0.397*** | story | 1.281*** | story | ||
| entity mentions | 0.285** | story | 0.006 | 1.085*** | story | |
| non-past tense | 0.947*** | non-story | – | – | 0.639 | |
| is comment (vs. post) | 0.612*** | non-story | n/a | n/a | n/a | |
| 2nd-person pronouns | 0.444*** | non-story | 0.285* | story | 0.017 | |
| sentence length | 0.259* | non-story | 0.012 | 0.828*** | non-story | |
| 1st-person plural pronouns | – | – | 0.106 | – | – | 0.154 |
| text length | – | – | 0.106 | 0.825*** | non-story | |